SUPERSIMPLE FIELDS AND DIVISION RINGS

ANAND PILLAY THOMAS SCANLON FRANK WAGNER

Abstract. It is proved that any supersimple field has trivial Brauer group, and more generally that any supersimple division ring is commutative. As prerequisites we prove several results about generic types in groups and fields whose theory is simple.